The fatality rate for lung cancer was 80 per 100’000 last year in East Europe and 16 deaths per 10’000 people in USA. Is it the truth that fatality rate is the same in Europe and USA?
Take significance level = 0,02
Can you please tell me how to solve that problem?
I've been looking for a easy solution but with no satisfying results, I still can't understand how to solve that.

Answer by jorel1380(3719)

Fatality rate for Europe:80/100000=.0008
Fatality rate for USA: 16/10000=.0016
Ratio of USA to Europe deaths:0.0016/0.0008=2:1
There were 2X as many fatalities in the USA than in Europe. ☺☺☺☺